# VASA-1: Lifelike Audio-Driven Talking Faces Generated in Real Time

### v1

@@ 5 Conclusion
In summary, our work presents VASA-1, an audio-driven talking face generation model that stands out for its efficient generation of realistic lip synchronization, vivid facial expressions, and naturalistic head movements from a single image and audio input. It significantly outperforms existing methods in delivering video quality and performance efficiency, demonstrating promising visual affective skills in the generated face videos. The technical cornerstone is an innovative holistic facial dynamics and head movement generation model that works in an expressive and disentangled face latent space.
@@
6 Social Impact and Responsible AI Considerations
Our research focuses on generating audio-driven visual affective skills for virtual AI avatars, aiming for positive applications. It is not intended to create content that is used to mislead or deceive. However, like other related content generation techniques, it could still potentially be misused for impersonating humans. We are opposed to any behavior to create misleading or harmful contents of real persons, and are interested in applying our technique for advancing forgery detection. Currently, the videos generated by this method still contain identifiable artifacts, and the numerical study shows that there’s still a gap to achieve the authenticity of real videos.
@@ Acknowledgments
We would like to thank our colleagues Zheng Zhang, Zhirong Wu, Shujie Liu, Dong Chen, Xu Tan and others for the valuable discussions and insightful suggestions for our project.

### v2

@@ 0 $\pm$1 $\pm$2 $\pm$3 $\pm$4
Figure 4: Ablation study on loss function $l_{consist}$ for disentangled latent space learning. We generate the results by only transferring the facial dynamics from source to target with head pose unchanged. $l_{consist}$ is crucial for decoupling subtle yet important facial dynamics from head pose.
Training data scale.
To validate the data scale influence and compare our model with previous methods at similar scales, we trained a diffusion model using only 10% of the data (i.e., 50K clips). As shown in Table 1, the model trained on this reduced dataset demonstrates comparable audio-lip and audio-pose synchronization to the full-dataset model, although the FVD and $\Delta$p metrics are not as good. Nonetheless, it still significantly outperforms previous methods across all metrics assessing synchronization, motion intensity, and video quality. This indicates that our approach remains highly effective even with much less data, and that increasing the dataset size enhances motion diversity.
Losses for latent space learning.
As described in Sec. 3.1, we introduce new losses $l_{consist}$ and $l_{cross\_id}$ to improve the disentanglement of facial dynamics, head pose, and face identity. To validate the effectiveness of $l_{consist}$, we transfer only facial dynamics from the source image to the target while keeping the target’s pose unchanged. Figure 4 shows that without $l_{consist}$, the latent model may struggle to replicate subtle facial dynamics such as side glances and lip asymmetries which are oftentimes coupled with head poses (e.g., a skewed mouth may coincide with a tilted head, and the gaze direction usually aligns with the head’s pose). Decoupling these subtle yet important dynamics are challenging without explicit constraints from $l_{consist}$.
We also evaluate the face identity loss $l_{cross\_id}$ for cross identity driving during training. We use all 108 subjects from the VoxCeleb2 test set for evaluation. For each subject, we chose the image that is closest to a frontal view from the first frames of all its clips to serve as the target image. Then we randomly selected 50 clips of other subjects as source videos, which leads to a total of 5,400 cross-reenactment clips. We calculate the facial identity preservation score by averaging the facial identity feature cosine similarity over all generated frames of all subjects. With the introduced face identity loss $l_{cross\_id}$, this identity preservation score of our results improved from $0.72$ to $0.80$.
@@ 5 Conclusion
In summary, our work presents an audio-driven talking face generation model that stands out for its efficient generation of realistic lip synchronization, vivid facial expressions, and naturalistic head movements from a single image and audio input. It significantly outperforms existing methods in delivering video quality and performance efficiency, demonstrating promising visual affective skills in the generated face videos. The technical cornerstone is an innovative holistic facial dynamics and head movement generation model that works in an expressive and disentangled face latent space.
@@ Acknowledgments
We would like to thank our colleagues Zheng Zhang, Zhirong Wu, Shujie Liu, Dong Chen, Xu Tan, Yu Deng, Lidong Zhou, and others for the valuable discussions and insightful suggestions for our project.
Appendix A Societal Impacts and Responsible AI Considerations
Our research focuses on generating audio-driven visual affective skills for virtual AI avatars, aiming for positive applications. It is not intended to create content that is used to mislead or deceive. However, like other related content generation techniques, it could still potentially be misused for impersonating humans. We are opposed to any behavior that creates misleading or harmful contents of real persons. Currently, the videos generated by this method still contain identifiable artifacts, and the numerical study in Section 4 shows that there’s still a gap to achieve the authenticity of real videos. Furthermore, we have trained a neural network based detector to distinguish real videos and those generated by our VASA-1, and the detector shows a $97.8\%$ accuracy for this task.
To combat potential misuse of our technique and other related ones and provide necessary safeguards, we are also working on applying our method for advancing face media forgery detection. Specifically, we are training generic face forgery detection models that incorporate our generated talking face videos as part of the training data. Our preliminary exploration shows that using our method to generate training data can lead to an obvious improvement of generality for the forgery detection models, and we’ll keep the community updated on new progresses.
Appendix B More Qualitative Evaluation, Comparison and Ablation Study
See Figure A.1 A.2 A.3 A.4 A.5 A.6 A.7.
Figure A.1: Disentanglement between identity and motion. In these examples, the same generated head and facial motion sequences are applied onto three different face images.
Figure A.2: Disentanglement between head pose and facial dynamics. From top to bottom: the raw generated sequence, applying generated poses with fixed initial facial dynamics, and applying generated facial dynamics with fixed initial head pose and pre-defined spinning poses, respectively.
Figure A.3: Generation results with out-of-distribution images (non-photorealistic) and audios (singing audios for the first two rows and non-English speech for the last row). Our method can still generate high quality videos well-aligned with the audios, although it was not trained on such data variations. See the supplementary video with audio for a better illustration of these results.
Figure A.4: Generation results from different methods with the input audio segment uttering “push ups”. See our supplementary video for a better illustration and comparison.
Figure A.5: Generation results from different methods with the input audio segment uttering “excruciating”. See our supplementary video for a better illustration and comparison.
Figure A.6: Generation results from different methods with the input audio segment uttering “what?”. See our supplementary video for a better illustration and comparison.
Figure A.7: Generation results from different methods with the input audio segment uttering “lots of questions”. See our supplementary video for a better illustration and comparison.